OpenStudy (anonymous):

Which value of x makes the following equation true? 2(x + 6) = x + 15 9 3 -3 15

OpenStudy (midhun.madhu1987):

2(x+6) = x+15 2x + 12 = x + 15 2x - x = 15-12 Find x and u get it...

OpenStudy (skullpatrol):

still confused?

OpenStudy (skullpatrol):